OpenStudy (anonymous):

Medal Awarded! If f(x)=1/x+2, what is f(3+h)? 1/3+h 1/h+5 1/(x+2)(3+h) (3+h)/(x+2)

OpenStudy (anonymous):

I'm guessing 1/h+5 but I'm not sure

OpenStudy (anonymous):

yep!

OpenStudy (ranga):

In f(x) is the entire (x+2) in the denominator or just the x?

OpenStudy (anonymous):

just x

OpenStudy (ranga):

If it just x, then f(3+h) = 1/(3+h) + 2 = 1/(3+h) + 2(3+h)/(3+h) = (2h+7)/(3+h) 2hich is not one of the choices.

OpenStudy (ranga):

Please use the Equation editor at the bottom of the reply box. Otherwise you will have to use parenthesis properly.

OpenStudy (anonymous):

.. u misread the question it's f(x) = 1/(x+2) not f(x) = 1/x + 2

OpenStudy (ranga):

That was exactly my earlier question: In f(x) is the entire (x+2) in the denominator or just the x?

OpenStudy (anonymous):

idk how you some how brought ln into this xD

OpenStudy (ranga):

There is a difference between: 1/x + 2 and 1/(x+2). With the equation editor they will respectively look like this: \[\Large \frac{ 1 }{ x } + 2 \quad ; \quad \frac{ 1 }{ x+2 }\] Thus when posting a question with the editor one should use parenthesis to make clear the whole thing is in the denominator.

OpenStudy (anonymous):

i just know how to do it with out ln :( too advanced!

OpenStudy (ranga):

That was not the natural logarithm ln (LN) but the word In (IN) in my earlier question.

OpenStudy (ranga):

I meant "when posting a question WITHOUT the editor one should use parenthesis" earlier.

OpenStudy (anonymous):

oh ok np :)